OpenStudy (anonymous):

Which is a counterexample to this conjecture? The sum of any two consecutive integers is a composite number.

OpenStudy (mertsj):

2+3=5

OpenStudy (anonymous):

not an asser

OpenStudy (anonymous):

looks like one to me

OpenStudy (anonymous):

A. 16 + 17 = 33 OR B. 10 + 11 = 21 OR C. 6 + 7 = 13 OR D. 7 + 8 = 15

OpenStudy (anonymous):

which one is not composite?

OpenStudy (anonymous):

no wich is

OpenStudy (kainui):

@raywolf1250 what is a composite number? What is a prime number? Can a number be both prime and composite at the same time?

OpenStudy (mertsj):

Consider 16+17=33 That is not the answer because 33 is a composite number since 33 = 3(11)

OpenStudy (mertsj):

Now consider the other answers and pick the sum whose answer is NOT composite. Therefore the answer is prime. Pick the one whose answer is prime.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

bye

OpenStudy (kainui):

@raywolf1250 we can help you solve your problem if you work with us, but we're not going to give you the answer.